julie's purse contains 32 coins consisting of nickels and quarters, worth a total of $3.60. How many quarters does she have?
Answer by Maths68(1474)

Let
Quarters=x
Nickels=32-x
Value of a Quarter = $0.25
Value of a Nickle = $0.05
Total value of all coins = $3.6
(Value of a Quarter)(Quarters)+(Value of a Nickel)(Nickels)= Total value of all coins
0.25x+0.05(32-x)=3.6
0.25x+0.05(32-x)=3.6
0.25x+1.6-0.05x=3.6
0.25x-0.05x=3.6-1.6
0.2x=3.6-1.6
0.2x=2
0.2x/0.2=2/0.2
x=10

Quarters=x=10
Nickels=32-x=32-10=22


Check
=========
0.25(10)+0.05(22)=3.6
2.5+1.1=3.6
3.6=3.6